# InstructMPC: A Human-LLM-in-the-Loop Framework for Context-Aware Control

### v2

@@ Abstract
1 Introduction
Model Predictive Control (MPC) is a versatile control strategy widely applied in domains such as energy management [1, 2], building control [3], and autonomous systems [4, 5, 6]. These applications leverage MPC’s ability to optimize system’s behavior over a predictive horizon, making it a cornerstone of modern control engineering.
@@ 1 Introduction
Language-based approaches harness natural language to personalize MPC systems for improved usability and flexibility. For instance, the LanguageMPC in [12] utilizes LLMs as decision-makers in autonomous driving, combining LLM reasoning with low-level controllers to enhance safety, efficiency, and generalizability in complex scenarios, though it does not support real-time human interaction. Meanwhile, the ChatMPC [13] processes natural language via BERT to update MPC specifications dynamically, reducing the data burden for personalization; however, it lacks state-of-the-art human-LLM interaction and theoretical performance guarantees.
@@
    Theoretical Guarantees: For linear dynamics, we provide theoretical performance (Theorem 4.1) guarantees for closed-loop LLM fine-tuning using Direct Preference Optimization (DPO). Unlike traditional online optimization methods that directly minimize regret, our approach decouples the DPO loss from the regret objective, which relies on MPC cost information and is challenging to compute online due to unknown future disturbances. In particular, Theorem 4.1 suggests a tailored loss function for LLM fine-tuning, achieving a regret bound of $O(\sqrt{T\log T})$ as derived in Corollary 4.1.
2 Preliminaries and Problem Formulation
Denote $[T]\coloneqq\{0,1,\ldots,T-1\}$. We consider a finite horizon linear dynamical system
$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad t\in[T]$ (1)
@@ 2 Preliminaries and Problem Formulation
There exists $K\in\mathbb{R}^{m\times n}$ such that $A-BK$ has a spectral radius $\rho$ less than $1$. [^2] Thus, the Gelfand’s formula implies that there exist $C>0$, $\rho\in(0,1)$ such that $\|A-BK\|^{t}\leq C\rho^{t}$ for all $t\geq 0$.
2.1 Model Predictive Control
@@
[…]
Only the first element of $u_{t:\mathcal{T}}^{\textsc{MPC}}$ is applied as a control input to the system, and the rests are discarded. However, in many real-world scenarios, the disturbances are hard to predict or measure [15, 16]. In this paper, we propose a novel framework that leverages contextual information to predict future disturbances and incorporates these predictions into the decision-making process.
2.2 Contextual MPC with Model Fine-Tuning
[…]
@@ 2.2 Contextual MPC with Model Fine-Tuning
where $J(\theta_{1},\ldots,\theta_{T})$ and $J(\theta^{\star})$ are the quadratic costs defined in (2a) induced by the MPC actions in (2.1) corresponding to the hyper-parameters $\theta_{1},\ldots,\theta_{T}$ and the optimal hyper-parameter $\theta^{\star}$, respectively. In other words, the regret in Equation 5 measures the difference between the cumulative loss incurred by $(\theta_{t}:t\in[T])$, and the cumulative loss incurred by the optimal $\theta^{\star}$ in hindsight. Next, we introduce the InstructMPC framework for contextual MPC.
3 InstructMPC
@@
Figure 1: System framework of InstructMPC. The blue lines represent interactions between InstructMPC and the environment where InstructMPC receives the state $x_{t}$, and outputs the control input $u_{t}$. The black lines represents the information loop, within which external contextual information is passed to the L2D module $g_{\theta}$ to produce predicted disturbances $\hat{w}_{t:\mathcal{T}|t}$. Then, the MPC controller $\pi_{\mathrm{MPC}}$ utilizes $\hat{w}_{t:\mathcal{T}|t}$ and the current state $x_{t}$ to determine a control input $u_{t}^{\mathrm{MPC}}$ via (2.1). After executing the MPC control input $u_{t}^{\mathrm{MPC}}$, the environment reveals the true disturbance $w_{t}$. The discrepancy between $w_{t}$ and $\hat{w}_{t}$ is then sent back to the L2D module as a loss signal.
3.1 Language to Distribution (L2D) Module
@@ 3.2 InstructMPC Framework
       Get predictions from $g_{\theta}$ with instruction $c_{t}$: $(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\leftarrow g_{\theta_{t}}(c_{t})$ Generate an action $u_{t}$ by solving MPC in (2.1) Update
@@
The L2D module is a neural network $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$, which is updated at every time $t$ through a loss function $L_{t-k+1}:\Theta\times\mathcal{W}^{2k}\rightarrow\mathbb{R}$, which also depends on disturbances $w_{t:\mathcal{T}}$ and predictions $\hat{w}_{t:\mathcal{T}|t}$. An example of such a loss function can be found in Corollary 4.1 in Section 4. At every time $t\in[T]$, the parameter $\theta$ of L2D module is updated according to:
@@ 4 Theoretical Guarantee
Note that the loss function used for updating $g_{\theta}$ may not be aligned with regret-optimal objective that requires the knowledge of future ground-truth predictions $(w_{t}:t\in[T])$ in (2.1). To quantify this mismatch, we define the loss discrepancy below.
@@
The loss discrepancy between two loss functions $L_{1}(\theta)$ and $L_{2}(\theta)$, $\textsf{LD}(L_{1},L_{2})$ is defined as
$\textsf{LD}(L_{1},L_{2})\coloneqq\sup_{\theta\in\Theta}\left\|\nabla_{\theta}L_{1}(\theta)-\nabla_{\theta}L_{2}(\theta)\right\|.$
The theorem below provides a regret bound (see Equation 5) for the proposed InstructMPC, which interacts with the dynamic system described via Equation 1.
Theorem 4.1.
@@
$+D\sum_{t=0}^{T-1}\textsf{LD}\left(L_{t},\psi_{t}^{\top}H\psi_{t}\right)+(k-1)GD,$
where we define
@@
By Theorem 3.2 in [20], given predictions $\hat{w}_{t:\mathcal{T}|t}$, the MPC solution $u_{t}^{\textsc{MPC}}$ to the problem defined in (2.1) is
@@
where $\psi_{t}(\theta)$ is defined in Theorem 4.1. For notational simplicity, we define $\psi_{t}\coloneqq\psi_{t}(\theta_{t})$ and $\phi_{t}\coloneqq\psi_{t}(\theta^{\star})$. By our model assumption, $Q,R\succ 0$ implies $H\succeq 0$, using the convexity of $x^{\top}Hx$,
@@
The bound in (7) is obtained by bounding the terms in Equation 9 separately, and the details are relegated to Appendix A. ∎
Furthermore, for a wide range of loss functions $(L_{t}:t\in[T])$, as examplified in Corollary 4.1 below, they exhibit a bounded discrepancy from the MPC cost with the discrepancy decaying exponentially as $k$ increases, i.e., $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k}$.
Corollary 4.1.
@@
there exist constants $C$ and $10<\rho<$ 1 such that for all $t\in[T]$,
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k},$
where $\psi_{t}(\theta)$ is defined in Theorem 4.1.
For the particular choice of $L_{t}$ in Corollary 4.1, the term $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ in the bound (7) of Theorem 4.1 simplifies to $CDT\rho^{k}$. This provides an insightful guideline for selecting $k$. Specifically, by setting $k=\log_{1/\rho}T,$ we can achieve a regret bound of $O(\sqrt{T\log T})$.
5 Numerical Experiment
In this section, we present a detailed demonstration of how the L2D module is fine-tuned in Subsection 5.1. We implement the L2D module with LLaMa-8B [21], a causal language model. Causal language models predict the next token in a sequence and return prediction scores for each vocabulary token, which aligns well with L2D’s requirement of generating scenario probabilities. To obtain these probabilities, we use the following prompt template: “[task description]. Given the current context: [context description], the expected scenario is [MASK]." For a predefined set of scenarios ($\mathcal{S}$ in Subsection 3.1), we evaluate the prediction score assigned by the model by inserting each candidate token into the [MASK] position. These scores are then normalized using a softmax function to produce a probability distribution.
In Subsection 5.2 and Subsection 5.3, we demonstrate the effectiveness of InstructMPC framework in two scenarios. The first is a robot tracking task, where an agent is required to follow an unknown trajectory. In this setting, the target position is revealed only immediately before the control decision is made. Moreover, environmental disturbances, such as wind force, are described textually and passed through the L2D module to assist the MPC controller. We plot the resulting trajectory to directly demonstrate the impact of contextual information. The second scenario is a real-world energy management task, in which human-provided instructions, such as news and scheduled appointments, are used to predict external disturbances. In this case, we further extend InstructMPC to operate on a nonlinear dynamical system, we also adopt a piecewise linear cost instead of the quadratic cost defined in eq. 2a, demonstrating the generalizability of InstructMPC in diverse system dynamics and different cost functions.
5.1 Fine-Tuning L2D Module
@@
5.2 Application 1: Robot Tracking
@@
$w_{t}\coloneqq Ay_{t}-y_{t+1}+Z_{t},$
$Z_{t}\coloneqq\begin{bmatrix}0\!\!\!&0\!\!\!&-0.2Z_{t}^{(1)}\!\!\!&-0.2Z_{t}^{(2)}\end{bmatrix}^{\top}\!\!\!,$
@@
To track the trajectory, the controller uses quadratic cost in Equation 2a with
@@
    Classic MPC: An MPC baseline using (2.1) where disturbance predictions are set to zero ($\hat{w}_{t:\mathcal{T}|t}=0$) for all $t$, equivalent to a Linear Quadratic Regulator (LQR) for the problem (2a)-(2b).
@@
Figure 2: Application 1: Robot Tracking (Section 5.2). Average robot tracking trajectories of three methods: 1. MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [21] model; 3. InstructMPC with LLaMA-8B [21] fine-tuned in a closed-loop. For each method, we ran $20$ episodes, and plot the average robot trajectory. The shadow region denotes the variance of trajectories.
5.3 Application 2: Building Energy Management
Problem description. We consider the problem of energy management for a campus building. The system consists of a battery for energy storage and a series of photovoltaic (PV) arrays to harvest the sunlight. Let $x_{t}\in[0,1]$ denote the state of charge (SoC) of the battery at time $t$, and $u_{t}\in[0,1]$ represent the charging rate, $u_{t}=1$ corresponds to charging at the maximum allowable power, and $u_{t}=0$ indicates no charging. The nonlinear system dynamics can be represented by: $x_{t+1}=f(x_{t},u_{t})+w_{t},$ where $f:\mathbb{R}^{2}\to\mathbb{R}$ captures the system’s inherent battery dynamics, $w_{t}\in\mathbb{R}$ represents the external disturbance, which can be attributed to the variations in solar irradiance or fluctuations in electricity demand. These disturbances are inferred from contextual information, including weather forecasts, human activity schedules, and natural language instructions. Examples of such contextual information include: “The CDC conference deadline is approaching, and many people are running experiments," or “I will be fine-tuning my LLaMA-1B model at 3 PM today, which will take approximately two hours." The objective of the energy management system is to minimize the total cost of purchasing electricity from the grid. This cost is computed using a piecewise linear function that reflects real-world peak and valley electricity prices.
Figure 3: Application 2: Building Energy Management (Section 5.3). Daily electricity cost comparison of three approaches: 1. MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [21] model; 3. InstructMPC with LLaMA-8B [21] fine-tuned in a closed-loop.

### v3

@@ Abstract
I Introduction
Model Predictive Control (MPC) is a versatile control strategy widely applied in domains such as energy management [1, 2, 3, 4], building control [5], and autonomous driving [6, 7]. These applications leverage MPC’s ability to optimize system’s behavior over a predictive horizon, making it a cornerstone of modern control engineering.
@@ I Introduction
Language-based approaches harness natural language to personalize MPC systems for improved usability and flexibility. For instance, the LanguageMPC in [13] utilizes LLMs as decision-makers in autonomous driving, combining LLM reasoning with low-level controllers to enhance safety, efficiency, and generalizability in complex scenarios, though it does not support real-time human interaction. Meanwhile, ChatMPC [14] uses BERT to classify natural language into a set of predefined circumstances, each linked to a predefined modification of MPC parameters; however, it is constrained by its predefined updating rule, and also lacks state-of-the-art human-LLM interaction and theoretical performance guarantees.
@@
    Theoretical Guarantees: For linear dynamics, we provide theoretical performance (Theorem IV.1) guarantees for closed-loop LLM fine-tuning using Direct Preference Optimization (DPO). Unlike traditional online optimization methods that directly minimize regret, our approach decouples the DPO loss from the regret objective, which relies on MPC cost information and is challenging to compute online due to unknown future disturbances. In particular, Theorem IV.1 suggests a tailored loss function for LLM fine-tuning, achieving a regret bound of $O(\sqrt{T\log T})$ as derived in Corollary IV.1.
II Preliminaries and Problem Formulation
Denote $T]\coloneqq\{0,1,\ldots,T-1\}$. We consider a finite horizon linear dynamical system
$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad t\in[T]\$(1)
@@ II Preliminaries and Problem Formulation
There exists $K\in\mathbb{R}^{m\times n}$ such that ABKA-BK has a spectral radius $\rho$ less than 11.222Thus, the Gelfand’s formula implies that there exist $C>0$, $\rho\in(0,1)$ such that $\|A-BK\|^{t}\leq C\rho^{t}$ for all $t\geq 0$.
II-A Model Predictive Control
@@
$\text{s.t. }x_{\tau+1}=Ax_{\tau}+Bu_{\tau}+\hat{w}_{\tau|t},\forall t\leq\tau<\mathcal{T}.$ (4)
Only the first element of $u_{t:\mathcal{T}}^{\textsc{MPC}}$ is applied as a control input to the system, and the rests are discarded. However, in many real-world scenarios, system disturbances are hard to predict [16, 17]. In this paper, we propose a novel framework that leverages contextual information to predict future disturbances and incorporates these predictions into the decision-making process.
II-B Contextual MPC with Model Fine-Tuning
Our goal is to improve the performance of the classic MPC defined in (II-A) by incorporating external contextual information at time $t\in[T]$, denoted by $c_{t}\in\mathcal{C}$ to predict future disturbances $\hat{w}_{t:\mathcal{T}|t}\in\mathcal{W}^{k}$. [^3] The predicted disturbance trajectory $\hat{w}_{t:\mathcal{T}|t}$ is generated in real-time, operating on the same time scale as the MPC decision-making process. Hence, the main challenge is to integrate a Neural Network (NN) $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$ to detokenize the contextual information into predictions for the downstream MPC, and refine these predictions by fine-tuning the model parameter $\theta\in\Theta$ after the real disturbances are revealed. Let $\|\cdot\|$ denote the $\ell_{2}$-norm of a vector. The hypothesis set $\Theta$ has bounded diameter so that $\|\theta-\theta^{\prime}\|\leq D$ for all $\theta,\theta^{\prime}\in\Theta$ and some constant $D>0$. We impose the following assumptions on the NN model $g_{\theta}$, which serve as approximations for widely used architectures, including transformers [18] and input convex neural networks [19].
@@ II-B Contextual MPC with Model Fine-Tuning
where $J(\theta_{1},\ldots,\theta_{T})$ and $J(\theta^{\star})$ are the quadratic costs defined in (2a) induced by the MPC actions in (II-A) corresponding to the hyper-parameters $\theta_{1},\ldots,\theta_{T}$ and the optimal hyper-parameter $\theta^{\star}$, respectively. In other words, the regret in Equation˜5 measures the difference between the cumulative loss incurred by $(\theta_{t}:t\in[T])$, and the cumulative loss incurred by the optimal $\theta^{\star}$ in hindsight. Next, we introduce the InstructMPC framework for contextual MPC.
III InstructMPC
@@
Figure 1: System framework of InstructMPC. The blue lines represent interactions between InstructMPC and the environment where InstructMPC receives the state $x_{t}$, and outputs the control input $u_{t}$. The black lines represents the information loop, within which external contextual information is passed to the L2D module $g_{\theta}$ to produce predicted disturbances $\hat{w}_{t:\mathcal{T}|t}$. Then, the MPC controller $\pi_{\mathrm{MPC}}$ utilizes $\hat{w}_{t:\mathcal{T}|t}$ and the current state $x_{t}$ to determine a control input $u_{t}^{\mathrm{MPC}}$ via (II-A). After executing the MPC control input $u_{t}^{\mathrm{MPC}}$, the environment reveals the true disturbance $w_{t}$. The discrepancy between $w_{t}$ and $\hat{w}_{t}$ is then sent back to the L2D module as a loss signal.
III-A Language to Distribution (L2D) Module
@@ III-B InstructMPC Framework
 Get predictions from $g_{\theta}$ with instruction $c_{t}$: $(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\leftarrow g_{\theta_{t}}(c_{t})$ Generate an action $u_{t}$ by solving MPC in (II-A) Update
@@
The L2D module is a neural network $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$, which is updated at every time $t$ through a loss function $L_{t-k+1}:\Theta\times\mathcal{W}^{2k}\rightarrow\mathbb{R}$, which also depends on disturbances $w_{t:\mathcal{T}}$ and predictions $\hat{w}_{t:\mathcal{T}|t}$. An example of such a loss function can be found in Corollary IV.1 in Section IV. At every time $t\in[T]$, the parameter $\theta$ of L2D module is updated according to:
@@ IV Theoretical Guarantee
A fundamental challenge is that the surrogate loss function used to train $g_{\theta}$ does not match the true, regret-optimal objective (II-A), which is unknown at training time. This misalignment creates the primary difficulty, i.e., ensuring that training the upstream L2D module $g_{\theta}$ actually improves the downstream MPC controller’s performance. Our solution is to define and analyze the loss discrepancy, a term that quantifies the divergence between the gradient of the L2D module surrogate loss $L$ and that of the true decision loss in Section˜II-A.
@@
The loss discrepancy between two loss functions $L_{1}(\theta)$ and $L_{2}(\theta)$, $\emph{{LD}}(L_{1},L_{2})$ is defined as
$\emph{{LD}}(L_{1},L_{2})\coloneqq\sup_{\theta\in\Theta}\left\|\nabla_{\theta}L_{1}(\theta)-\nabla_{\theta}L_{2}(\theta)\right\|.$
The theorem below provides a regret bound (see Equation˜5) for the proposed InstructMPC, which interacts with the dynamic system described via Equation˜1.
Theorem IV.1.
@@
$+D\sum_{t=0}^{T-1}\emph{{LD}}\left(L_{t},\psi_{t}^{\top}H\psi_{t}\right)+(k-1)GD,$
where $H$ is defined in Lemma 1, and we define
@@
By Theorem 3.2 in [21], given predictions $\hat{w}_{t:\mathcal{T}|t}$, the MPC solution $u_{t}^{\textsc{MPC}}$ to the problem defined in (II-A) is
@@
where $\psi_{t}(\theta)$ is defined in Theorem IV.1. For notational simplicity, we define $\psi_{t}\coloneqq\psi_{t}(\theta_{t})$ and $\phi_{t}\coloneqq\psi_{t}(\theta^{\star})$. By our model assumption, $Q,R\succ 0$ implies $H\succeq 0$, using the convexity of $x^{\top}Hx$,
@@
The bound in (7) is obtained by bounding the terms in Equation˜9 separately, and the details are relegated to Appendix Appendix A: Proof of Theorem IV.1. ∎
Furthermore, for a wide range of loss functions $(L_{t}:t\in[T])$, as examplified in Corollary IV.1 below, they exhibit a bounded discrepancy from the MPC cost with the discrepancy decaying exponentially as $k$ increases, i.e., $\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k}$.
Corollary IV.1.
@@
there exist constants $C$ and $0<\rho<1$ such that
$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k},\ \text{ for all }t\in[T],$
where $\psi_{t}(\theta)$ is defined in Theorem IV.1.
For the particular choice of $L_{t}$ in Corollary IV.1, the term $D\sum_{t=0}^{T-1}{\texttt{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})$ in the bound (7) of Theorem IV.1 simplifies to $CDT\rho^{k}$. This provides an insightful guideline for selecting $k$. Specifically, by setting $k=\log_{1/\rho}T,$ we can achieve a regret bound of $O(\sqrt{T\log T})$.
V Numerical Experiment
In this section, we present a detailed demonstration of how the L2D module is fine-tuned in Subsection V-A. We implement the L2D module with LLaMa-8B [22], a causal language model. Causal language models predict the next token in a sequence and return prediction scores for each vocabulary token, which aligns well with L2D’s requirement of generating scenario probabilities. To obtain these probabilities, we use the following prompt template: “[task description]. Given the current context: [context description], the expected scenario is [MASK].” For a predefined set of scenarios ($\mathcal{S}$ in Subsection III-A), we evaluate the prediction score assigned by the model by inserting each candidate token into the [MASK] position. These scores are then normalized using a softmax function to produce a probability distribution.
In Subsection V-B and Subsection V-C, we demonstrate the effectiveness of InstructMPC framework in two scenarios. The first is a robot tracking task, where an agent is required to follow an unknown trajectory. In this setting, the target position is revealed only immediately before the control decision is made. Moreover, environmental disturbances, such as wind force, are described textually and passed through the L2D module to assist the MPC controller. We plot the resulting trajectory to directly demonstrate the impact of contextual information. The second scenario is a real-world energy management task, in which human-provided instructions, such as news and scheduled appointments, are used to predict external disturbances. In this case, we further extend InstructMPC to operate on a nonlinear dynamical system, we also adopt a piecewise linear cost instead of the quadratic cost defined in eq.˜2a, demonstrating the generalizability of InstructMPC in diverse system dynamics and different cost functions.
V-A Fine-Tuning L2D Module
@@
V-B Application 1: Robot Tracking
@@
$w_{t}\coloneqq Ay_{t}-y_{t+1}+Z_{t},Z_{t}\coloneqq\begin{bmatrix}0\!\!\!&0\!\!\!&-0.2Z_{t}^{(1)}\!\!\!&-0.2Z_{t}^{(2)}\end{bmatrix}^{\top}\!\!\!$,
@@
To track the trajectory, the controller uses quadratic cost in Equation˜2a with
@@
    Classic MPC: An MPC baseline using (II-A) where disturbance predictions are set to zero ($\hat{w}_{t:\mathcal{T}|t}=0$) for all $t$, equivalent to a Linear Quadratic Regulator (LQR) for the problem (2a)-(2b).
@@
Figure 2: Application 1: Robot Tracking (Section˜V-B). Average robot tracking trajectories of three methods: 1. MPC in (II-A) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [22] model; 3. InstructMPC with LLaMA-8B [22] fine-tuned in a closed-loop. For each method, we ran 2020 episodes, and plot the average robot trajectory. The shadow region denotes the variance of trajectories.
V-C Application 2: Building Energy Management
Problem description. We consider the problem of energy management for a campus building. The system consists of a battery for energy storage and a series of photovoltaic (PV) arrays to harvest the sunlight. Let $x_{t}\in[0,1]$ denote the state of charge (SoC) of the battery at time $t$, and $u_{t}\in[0,1]$ represent the charging rate, $u_{t}=1$ corresponds to charging at the maximum allowable power, and $u_{t}=0$ indicates no charging. The nonlinear system dynamics can be represented by: $x_{t+1}=f(x_{t},u_{t})+w_{t},$ where $f:\mathbb{R}^{2}\to\mathbb{R}$ captures the system’s inherent battery dynamics, $w_{t}\in\mathbb{R}$ represents the external disturbance, which can be attributed to the variations in solar irradiance or fluctuations in electricity demand. These disturbances are inferred from contextual information, including weather forecasts, human activity schedules, and natural language instructions. Examples of such contextual information include: “The CDC conference deadline is approaching, and many people are running experiments,” or “I will be fine-tuning my LLaMA-1B model at 3 PM today, which will take approximately two hours.” The objective of the energy management system is to minimize the total cost of purchasing electricity from the grid. This cost is computed using a piecewise linear function that reflects real-world peak and valley electricity prices.
Figure 3: Application 2: Building Energy Management (Section˜V-C). Daily electricity cost comparison of three approaches: 1. MPC in (II-A) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [22] model; 3. InstructMPC with LLaMA-8B [22] fine-tuned in a closed-loop.